### v1

# A Practical Approach Towards Inertia Estimation Using Ambient Synchrophasor Data

Abstract
Real-time tracking of inertia is important because it reflects the power system’s ability to withstand contingencies and maintain frequency security. This paper proposes a practical approach to estimate inertia using ambient phasor measurement unit (PMU) data and a partitioned form of the swing equation. The approach accounts for (bounded) uncertainties in network parameters and PMU measurements, enabling precise estimation of inertia and damping constants, as well as mechanical power inputs. Instead of assuming constant mechanical power input throughout, the approach leverages knowledge of power system operations to determine intervals when it is actually constant to maintain estimation consistency. Simulation results on the IEEE 14-bus system and IEEE 39-bus system integrated with renewable energy sources affirm the method’s accuracy and applicability.
Index Terms:
Ambient data, Bounded data uncertainty, Frequency divider, Inertia estimation, Phasor measurement unit
I Introduction
With spatially-dispersed converter-interfaced generation (CIG) gradually replacing centrally-located synchronous generation, it has become necessary to track system inertia in real-time [1]. Real-time methods for inertia estimation primarily comprise phasor measurement unit (PMU) data-driven approaches, and can be disturbance-based, perturbation-based, or ambient data-based [2]. The dependence on events and judicious selection of the probing signal limit the practical use of the disturbance-based and perturbation-based approaches, respectively [3]. The ambient-data based inertia estimation approaches often assume constant mechanical power input [4] and ignore uncertainties in the network parameters. This paper advances the state-of-the-art by performing real-time inertia estimation using ambient PMU data while accounting for the variations in mechanical power inputs and uncertainties in network parameters as well as synchrophasor measurements.
We first estimate the rotor speeds and acceleration from PMUs placed at a few of the network buses by leveraging the frequency divider formula [5], and then solve the resulting linear estimation problem in the presence of bounded uncertainties. Next, we determine the inertia constant ($H$) and damping constant ($D$) of individual generators from the swing equation while treating the mechanical power input ($p_{m}$) as a constant. However, since we do not know beforehand the length of the actual windows for which $p_{m}$ is constant, considerable fluctuations occur in the $H$ and $D$ estimates. We use the fluctuations in the estimates to figure out time intervals for which $p_{m}$ is an unknown constant, and then solve for all three unknown parameters ($H$, $D$, $p_{m}$) through a partitioned form of the swing equation. We draw insights from actual power system operations to ensure practicality of the proposed approach. Our main contributions are:
  1. 1.
    Along with inertia, we also determine damping constant and mechanical power input for the windows for which the mechanical power input is actually constant.
  2. 2.
    We demonstrate that variations in mechanical power inputs can be inferred from fluctuations in inertia and damping estimates leading to a relaxation of the assumption of constant mechanical power input throughout.
  3. 3.
    We exploit the bounded nature of the perturbations in network parameters and PMU data to treat the estimation problems as bounded data uncertainty (BDU) problems.
II Proposed Approach for Rotor Speed and Acceleration Estimation
The frequency divider formula introduced in [5] establishes a linear relationship between rotor speeds of synchronous machines and the frequency of network buses. Combining this formula with the knowledge of network topology and bus frequencies obtained via PMUs, one can estimate the rotor speeds (and subsequently, acceleration), which can then be used to estimate inertia from the swing equation. However, network parameters are subject to changes and PMU measurements have errors, implying that the estimation problem will be subject to uncertainties. These aspects of the problem are discussed in this section, while the proposed approach for inertia estimation is described in the next section.
II-A Frequency Divider and Spatial Correlation
Consider a power system having $N$ buses and $M$ point of interconnection (POI) buses to which generators are connected. For such a system, the frequency divider formula can be mathematically expressed in the following way [5]:
$0_{N}=B_{BG}\omega_{G}(t)+B_{BB}\omega_{B}(t),$ (1)
where $\omega_{B}(t)\in\mathbb{R}^{N}$ is the vector of bus frequencies; $\omega_{G}(t)\in\mathbb{R}^{M}$ is the vector of rotor speeds; and the matrices are:
  • $B_{BG}\in\mathbb{R}^{N\times M}$ denotes the susceptance between the generators and the buses to which they are connected,
  • $B_{BB}=B_{BUS}+B_{GG}\in\mathbb{R}^{N\times N}$,
  • $B_{BUS}\in\mathbb{R}^{N\times N}$ is the imaginary part of the complex-valued network admittance matrix,
  • $B_{GG}\in\mathbb{R}^{N\times N}$ is a diagonal matrix with $B_{GG}(i,i)\neq 0$ if a generator is present at bus $i$, and its value is the inverse of the generator’s internal reactance.
For ease of notation, we drop the time argument $t$ from both $\omega_{G}(t)$ and $\omega_{B}(t)$. Equation (1) can now be partitioned as:
$\begin{bmatrix}0_{M}\\ 0_{N}\end{bmatrix}=\begin{bmatrix}\operatorname{diag}(b_{Gb})&B_{Bb}\\ 0_{N\times M}&B_{Bl}\end{bmatrix}\begin{bmatrix}\omega_{G}\\ \omega_{B}\end{bmatrix}$ (2)
where $\operatorname{diag}(b_{Gb})\in\mathbb{R}^{M\times M}$ is a diagonal matrix whose entries are the inverse of the combined reactances of the generator’s internal reactance and the reactance of the line and transformer located between the generator and the POI, $B_{Bb}$ captures the relation between frequency of POI buses and rotor speeds, and $B_{Bl}$ relates the frequency of one bus with that of electrically connected buses. From the first row of (2), we get:
${\omega}_{G}=-\operatorname{diag}(b_{Gb})^{-1}B_{Bb}{\omega}_{B}$ (3)
If PMUs measure $L$ ($L\leq N$) of the $N$ bus frequencies, then exploiting the frequency spatial correlations that exist in the power system, we can write:
$\omega_{B,me}=E{\omega}_{B}$ (4)
where $\omega_{B,me}\in\mathbb{R}^{L\times 1}$ denotes the measured bus frequencies, and $E$ $\in\mathbb{R}^{L\times N}$ is an identity-like matrix whose rows corresponding to bus frequencies that are not measured by PMUs, have been removed. Next, by combining (4) with the second row of (2), the following relation between the generators’ rotor speeds and measured bus frequencies can be established [6]:
${\omega}_{G}$ =$=\left\{\operatorname{diag}(b_{Gb})^{-1}B_{Bb}\begin{bmatrix}E\\ B_{Bl}\end{bmatrix}^{+}\begin{bmatrix}I_{L}\\ 0_{N\times L}\end{bmatrix}\right\}\omega_{B,me}$ (5)
where, $I_{L}$ $\in\mathbb{R}^{L\times L}$ is the identity matrix, and $[\cdot]^{+}$ denotes the pseudo-inverse. Finally, by denoting the terms in curly brackets in (5) by $C\in\mathbb{R}^{M\times L}$, we establish the following linear relation between rotor speeds and measured bus frequencies:
${\omega}_{G}=C\omega_{B,me}$ (6)
To uniquely estimate all rotor speeds using (6), at a minimum, every POI bus frequency must be measured by PMUs. Going forward, we assume that is the case and set $L=M$. Note that one can show $C$ is full rank in this case. Finally, since PMUs give rate-of-change-of-frequency measurements and $C$ is time-invariant (as long as topology does not change), (6) can be extended to solve for rotor acceleration as well.
II-B Estimation in Presence of Bounded Data Uncertainties
It has been documented in prior literature that network parameter uncertainties are bounded to within $\pm$30of their database values [7], while errors in the PMU-measured bus frequencies in the U.S. lie within $\pm$ 0.008 Hz [8]. To effectively incorporate knowledge of these bounds into our problem formulation, we treat the estimation problems as BDU problems. Taking (6) as an example of a BDU problem, we solve for $\hat{\omega}_{G}$ by performing the following $\operatorname{min-max}$ optimization:
$\min_{{\omega}_{G}}\max_{\begin{array}[]{c}\scriptstyle\|\delta\mathbb{A}\|_{2}\leq\eta\\[-4.0pt] \scriptstyle\|\delta{\omega}_{B,me}\|_{2}\leq\eta_{b}\\[-4.0pt] \end{array}}{\|(\mathbb{A}+\delta\mathbf{\mathbb{A}}){\omega}_{G}-({\omega}_{B,me}+\delta{\omega}_{B,me})\|_{2}}$
where $\mathbb{A}=C^{-1}$, $\delta\mathbf{\mathbb{A}}$ and $\delta{\omega}_{B,me}$ are unknown perturbations in $\mathbb{A}$ and ${\omega}_{B,me}$, and $\eta$ and $\eta_{b}$ are known bounds on the perturbations. The solution, $\hat{\omega}_{G}$, is constructed as follows [9]:
  • Perform singular value decomposition (SVD) of $\mathbb{A}$ and obtain unitary matrices $U\in\mathbb{R}^{M\times M}$ and $V\in\mathbb{R}^{M\times M}$ which represent rotations in space, and diagonal matrix $S\in\mathbb{R}^{M\times M}$ which scales each coordinate by the singular value of $\mathbb{A}$. Define $w_{B,me}=U^{\top}{\omega}_{B,me}$.
  • Introduce the secular function: $\mathbb{G(\psi)}={w_{B,me}}^{\top}(S^{2}-\eta^{2}I)(S^{2}+\psi I)^{-2}$, where $\psi$ is a positive solution of
    $\psi=\frac{\eta\sqrt{\psi^{2}\|(S^{2}+\psi I_{M})^{-1}{{w}_{B,me}}\|_{2}}}{\|S(S^{2}+\psi I_{M})^{-1}{{w}_{B,me}}\|_{2}}$ (7)
    where $I_{M}\in R^{M\times M}$ is the identity matrix.
  • Define
    $\tau_{1}=\frac{\|S^{-1}{w_{B,me}}\|_{2}}{\|S^{-2}{w_{B,me}}\|_{2}}\>\>\>\mathrm{and}\>\>\tau_{2}=\frac{\|\mathbb{A}^{\top}{\omega_{B,me}}\|_{2}}{\|{\omega_{B,me}}\|_{2}}$
    With these definitions, the solution $\hat{\omega}_{G}$ is given by [9]:
    1. 1.
      For $\eta\geq\tau_{2}$, the unique solution is $\hat{\omega}_{G}=0$.
    2. 2.
      For $\tau_{1}<\eta<\tau_{2}$, the unique solution is given by $\hat{\omega}_{G}=(\mathbb{A}^{\top}\mathbb{A}+\hat{\psi}I)^{-1}\mathbb{A}^{\top}{\omega_{B,me}}$, where $\hat{\psi}$ is the positive root of the secular equation, $\mathbb{G}(\hat{\psi})$ = 0.
    3. 3.
      For $\eta\leq\tau_{1}$, the unique solution is $\hat{\omega}_{G}=VS^{-1}w_{B,me}=\mathbb{A}^{+}{\omega_{B,me}}$.
    4. 4.
      For $\eta=\tau_{1}=\tau_{2}$, there are infinitely many solutions: $\hat{\omega}_{G}=\beta VS^{-1}w_{B,me}=\beta\mathbb{A}^{+}{\omega_{B,me}}$, for any $0\leq\beta\leq 1$.
The above-mentioned steps are followed for solving all the BDU problems encountered in this paper.
III Proposed Approach for Inertia, Damping, and Mechanical Power Input Estimation
The swing equation for the $j$-th generator is:
$2H_{j}\dot{\omega}_{G,j}(t)+D_{j}{\omega_{G,j}(t)-p_{m,j}(t)=-p_{e,j}(t)}$ (8)
where $H_{j}$ denotes its inertia constant, $D_{j}$ denotes its damping constant, $\dot{\omega}_{G,j}$ denotes its rotor acceleration, $\omega_{G,j}$ denotes its rotor speed, $p_{m,j}$ denotes its mechanical power input, and $p_{e,j}$ denotes its electrical power output. Dropping the subscript $j$ and focusing on a single generator for $t=0,\ldots,N$, we get:
$\underbrace{\begin{bmatrix}\dot{\omega_{G}}(0)&\omega_{G}(0)&-1\\ \dot{\omega_{G}}(1)&\omega_{G}(1)&-1\\ \vdots&\vdots&\vdots\\ \dot{\omega_{G}}(N)&\omega_{G}(N)&-1\end{bmatrix}}_{A}\underbrace{\begin{bmatrix}\bar{M}\\ D\\ \bar{p}_{m}\end{bmatrix}}_{x}$ $=\underbrace{\begin{bmatrix}\bar{p}_{e}(0)\\ \bar{p}_{e}(1)\\ \vdots\\ \bar{p}_{e}(N)\end{bmatrix}}_{b}$ (9)
where $\bar{M}=2H$, $\bar{p}_{m}=p_{m}$, and $\bar{p}_{e}(t)=-p_{e}(t)$. It is clear that (9) will give reasonable estimates for the three unknown parameters in $x$ only when $p_{m}$ is a constant. We now leverage our knowledge of power system operations to better understand the conditions under which $p_{m}$ varies.
III-A Behavior of Mechanical Power Input
The mechanical power input of a generator usually changes in response to the power system’s economic dispatch. Furthermore, even if the economic dispatch commands coming from the independent system operator have a longer interval (say, $5$ minutes), a utility may want to change their generation at shorter intervals (say, $\mathrm{tens}$ of seconds), particularly if the system has a large number of CIGs whose inputs are a function of weather conditions. Thus, a typical variation in the $p_{m}$ of a generator over an economic dispatch period ($T$) may look similar to the plot shown in Fig. 1. In the figure, $p_{m}$ ramps considerably in the intervals $d_{1}$, $d_{2}$, and $d_{3}$, and is relatively flat in other time intervals. However, even though it may be flat, its value is not the same. Additionally, one cannot measure $p_{m}$ in real-time using a PMU. Therefore, even if $p_{m}$ remains constant, one cannot know its value by direct measurement.
Figure 1: Realistic variations in mechanical power input
It is also clear from Fig. 1 that $p_{m}$ can change significantly between the start and end of an economic dispatch period. Lastly, because $p_{m}$ can change within the economic dispatch period, the durations for which $p_{m}$ is constant (relatively flat in Fig. 1) might not be known in advance. These practical considerations explain why (9) is unsuitable for inertia estimation when $p_{m}$ changes. A strategy to circumvent this problem is presented below.
III-B A Partitioned Formulation for Inertia Estimation
Since PMUs produce outputs at a much faster rate than the frequency with which $p_{m}$ changes, one can create rolling windows of duration much smaller than the economic dispatch period ($T$ in Fig. 1), and solve (9) for each window. The outcome will be that the estimates over consecutive windows will become inconsistent whenever $d_{1}$, $d_{2}$, $d_{3}$ appear inside the windows. However, the inconsistency of the estimates will reveal the durations over which $p_{m}$ is varying. This information is leveraged below to create a partitioned formulation of the swing equation for inertia estimation.
Using the knowledge of when the estimates obtained from (9) have become inconsistent within the economic dispatch period, we partition the discrete-time interval $t=0,\ldots,T$ into $k$ non-overlapping intervals, denoted as $[t_{i-1}+d_{i-1},t_{i})$ for $i=1,\ldots,k$, where $t_{0}=0$, $d_{0}=0$, and $t_{k}=T$. Note that each of these intervals can have an arbitrary length. Now, since $p_{m}$ remains constant within each interval but may differ between intervals, we have $\bar{p}_{m}(t)=\bar{p}_{m_{i}}$ for $t\in[t_{i-1}+d_{i-1},t_{i})$. Finally, for $i\in\{1,\ldots,k\}$, it follows that:
$\begin{bmatrix}\bar{p}_{m}(t_{0}+d_{0})\\ \vdots\\ \bar{p}_{m}(t_{1})\\ \hline\cr\bar{p}_{m}(t_{1}+d_{1})\\ \vdots\\ \bar{p}_{m}(t_{2})\\ \hline\cr\vdots\\ \hline\cr\bar{p}_{m}(t_{k-1}+d_{k-1})\\ \vdots\\ \bar{p}_{m}(t_{k})\\ \end{bmatrix}$ $=\underbrace{\begin{bmatrix}1&0&0&\ldots&0\\ 1&0&0&\ldots&0\\ \vdots\\ 1&0&0&\ldots&0\\ \hline\cr 0&1&0&\ldots&0\\ 0&1&0&\ldots&0\\ \vdots\\ 0&1&0&\ldots&0\\ \hline\cr\vdots&\vdots&\vdots&\vdots\\ \hline\cr 0&0&0&\ldots&1\\ 0&0&0&\ldots&1\\ \vdots\\ 0&0&0&\ldots&1\\ \end{bmatrix}}_{A_{2}}\begin{bmatrix}\bar{p}_{m_{1}}\\ \bar{p}_{m_{2}}\\ \vdots\\ \bar{p}_{m_{k}}\end{bmatrix}$ (10)
Based on the partitioned model in (10), we define $\boldsymbol{\omega}_{G}(t_{i})=[\omega_{G}(t_{i-1}+d_{i-1}),\ldots,\omega_{G}(t_{i})]^{\top}$ for $i\in\{1,\ldots,k\}$. Similarly, we also define $\dot{\boldsymbol{\omega}}_{G}(t_{i})$ and $\bar{\boldsymbol{p}}_{e}(t_{i})$. Combining these definitions and (10) with (9), we get:
$\underbrace{\begin{bmatrix}\dot{\boldsymbol{\omega}}_{G}(t_{1})&\boldsymbol{\omega}_{G}(t_{1})\\ \dot{\boldsymbol{\omega}}_{G}(t_{2})&\boldsymbol{\omega}_{G}(t_{1})\\ \vdots&\vdots\\ \dot{\boldsymbol{\omega}}_{G}(t_{k})&\boldsymbol{\omega}_{G}(t_{k})\end{bmatrix}}_{A_{1}}\begin{bmatrix}\bar{M}\\ D\\ \end{bmatrix}-A_{2}\begin{bmatrix}\bar{p}_{m_{1}}\\ \bar{p}_{m_{2}}\\ \vdots\\ \bar{p}_{m_{k}}\end{bmatrix}$ $=\underbrace{\begin{bmatrix}\bar{\boldsymbol{p}}_{e}(t_{1})\\ \bar{\boldsymbol{p}}_{e}(t_{2})\\ \vdots\\ \bar{\boldsymbol{p}}_{e}(t_{k})\end{bmatrix}}_{b_{1}}$ (11)
Finally, by rearranging the terms in (11), we get:
$\begin{bmatrix}A_{1}&|&-A_{2}\end{bmatrix}\begin{bmatrix}\bar{M}\\ D\\ \hline\cr\bar{p}_{m_{1}}\\ \vdots\\ \bar{p}_{m_{k}}\end{bmatrix}=b_{1}$ (12)
where the number of unknown parameters are $k+2$. Therefore, as long as the number of rows of $A_{1}$, $A_{2}$, and $b_{1}$ are greater than $k+2$ (which is always the case since PMUs produce outputs faster than the frequency with which $p_{m}$ changes), one can uniquely estimate these parameters.
Now, since $A_{1}$ is derived using the methodology described in Section II-A and $b_{1}$ is obtained from PMU measurements, $A_{1}$ and $b_{1}$ can have bounded uncertainties. Therefore, (12) will be treated as a BDU problem [9] and solved using the methodology described in Section II-B. Finally, the above-mentioned procedure will be repeated for all the generators present in the system to calculate the system-level inertia constant, $H_{sys}$, using the following relation:
$\hat{H}_{sys}=\frac{\sum_{j}\hat{H}_{j}S_{j}}{\sum_{j}S_{j}}$ (13)
where, $S_{j}$ denotes the rated capacity of generator $j$, and $\hat{H}_{j}$ is the estimated inertia constant of generator $j$ obtained by solving (12). Note that the proposed approach estimates system-level inertia constant for the immediately prior economic dispatch period. This is not a problem during actual implementation because inertia changes slowly and a delay of one economic dispatch period for getting accurate and consistent inertia estimates is an acceptable compromise.
IV Simulation Results
We evaluate the proposed approach on the IEEE 14-bus and 39-bus systems. The former has no CIGs, but the latter has three CIGs at buses 2, 29, and 39. We generate data by creating random fluctuations in the loads by varying them within ±$\pm 0.1\%$ of their actual value [10]. Bounded perturbations specified in [7] and [8] are added to the network parameters and PMU measurements, respectively. The $p_{e}$ of the generators is changed every $8$ seconds, and the dynamic simulations are run for $40$ seconds in PSS/E. We use the absolute relative error ($\mathrm{ARE}$) metric for evaluating the performance.
IV-A IEEE 14-Bus System
This system has five generators labeled {$G_{1}$, $G_{2}$, $G_{3}$, $G_{4}$, $G_{5}$}. The variation in $p_{e}$ and $p_{m}$ for $G_{1}$ is shown in Fig. 2. However, the knowledge of the interval after which the change occurs ($8$ seconds, in this case), is not known a priori. Therefore, we first solve (9) over the entire $40$-second time interval using a rolling window of size $0.33$ seconds. Figs. 3 and 4 demonstrate that the fluctuations in $p_{m}$ (shown in Fig. 2) introduce extremely high variability in the estimates of $H$ and $D$. Note that the regions of the plots of %$\%H_{\mathrm{ARE}}$ and %$\%D_{\mathrm{ARE}}$ which are blank (e.g., first $120$ rolling windows in Fig. 4) were due to the estimation algorithm not converging because of the ill-conditioning of the linear system of equations. One way to avoid this problem is by increasing the size of the rolling window. The optimal window size should be such that (i) the linear estimation problems derived in Section III are rank-sufficient, and (ii) there are multiple windows in which $p_{m}$ has the time to settle down and become flat.
Figure 2: $p_{e}$ vs. $p_{m}$ (in p.u.) on the same base for $G_{1}$
[…]
Now, by analyzing the variability in the estimates of $H$ and $D$, we identify time intervals where $p_{m}$ remains constant. Then, we use (12) to estimate $H$, $D$, and $p_{m}$ for those time intervals only. The results are shown in Table I for all five generators of the 14-bus system. Note that the entries in the last column indicate the average %$\%\mathrm{ARE}$ for $p_{m}$ across different time intervals. The very low values of the %$\%\mathrm{ARE}$ confirm the robustness of the proposed approach.
[…]
H D $p_{m,avg}$
1 0.0115 0.41 0.6
2 0.0322 0.24 0.3
3 0.0198 0.21 0.09
4 0.0272 0.17 0.4
5 0.0121 0.20 0.5
IV-B IEEE 39-Bus System with Renewable Energy Resources
The modified IEEE 39-bus system consists of 3 renewable resources and 10 generators labeled {$G_{1}$, $\ldots$, $G_{10}$}. Estimation of $H$ and $D$ using (9) over the $40$-second interval results in Figs. 5 and 6. The figures confirm that the estimated values of $H$ and $D$ fluctuate considerably. Note that in contrast to Figs. 3 and 4 which showed the fluctuations in the %$\mathrm{ARE}$, Figs. 5 and 6 depict the variations in the actual estimates. However, the inferences drawn are the same (namely, that changes in $p_{m}$ severely deteriorate estimation performance). By examining these plots, intervals in which $p_{m}$ remains constant are identified and the remaining rolling windows are discarded. The %$\%\mathrm{ARE}$ of $H$, $D$, and $p_{m}$ estimates obtained using (12), after omitting the intervals where $p_{m}$ is varying, are shown in Table II. The %$\%\mathrm{ARE}$ values (particularly for $\hat{H}$) are even lower for this system confirming the accuracy and widespread applicability of the proposed approach.
Figure 5: Fluctuations in $\hat{H}$ for $G_{3}$ for varying mechanical power input
Figure 6: Fluctuations in $D$ estimates for $G_{3}$ in presence of varying mechanical power input
[…]
H ×$\times 10^{-4}$ D $p_{m,avg}$
1 3.6592 0.0853 0.0292
2 3.4466 1.51 0.0105
3 3.6608 0.15 0.0405
4 1.1032 0.11 0.0014
5 4.9260 0.31 0.0441
6 2.7518 0.26 0.0889
7 2.1293 0.23 0.0774
8 2.4934 0.93 0.0358
9 4.7090 1.66 0.0808
10 4.2717 1.49 0.0784
Lastly, Table III presents the system-level inertia estimates for the IEEE 14-bus and 39-bus systems. As seen from the table, the calculated system-level inertia constant values are very close to the true values (obtained from PSS/E dynamic data for the two test systems), demonstrating the effectiveness of the proposed inertia estimation method in presence of bounded uncertainties in the network parameters and PMU measurements and changing mechanical power inputs.
TABLE III: Estimated system-level inertia constant ($\hat{H}_{sys}$) using partitioned form of swing equation
Test System True value of $\hat{H}_{sys}$ Estimated value of $\hat{H}_{sys}$
14-bus 4.9400 seconds 4.9404 seconds
39-bus 3.2000 seconds 3.1999 seconds
V Conclusion
This paper presents a novel approach for estimating three parameters—namely, inertia constant, damping constant, and mechanical power input—of every generator from ambient PMU data. The proposed approach relaxes the constant mechanical power input assumption and accounts for the presence of bounded uncertainties in the network parameters and PMU measurements. First, rotor speed and acceleration are estimated for all the generators from PMUs placed only at the POI buses by embedding frequency spatial correlations into the frequency divider formula. Then, these estimates are inserted into the swing equation to check for consistency of the inertia and damping estimates. Considerable variations in the resulting estimates are indicators of changes in mechanical power inputs. This knowledge is used to determine intervals for which the mechanical power input does not change. Finally, these intervals are fed into a partitioned form of the swing equation to estimate the three desired parameters as well as the system-level inertia constant. The results obtained using two IEEE test systems affirm the accuracy, robustness, and widespread applicability of the proposed approach. The focus of future research will be on investigating this approach in the presence of intermittency and variability typically encountered by CIGs in the field.